# Cross-Platform Benchmarking of Near‑Term Quantum Optimisation Algorithms

### v4

@@ I Introduction
To perform our cross-platform benchmarking we chose the problem of finding the lowest-energy configuration of defective graphene structures, based on the formulation found in Ref. [28]. We use the graphene defect problem solely as a representative, high density instance on which to test our benchmarking framework. Suppose we are tasked with removing a specific number of atoms from a graphene sheet of $N$ sites. The goal is to find which atoms, when removed, will result in a configuration with the lowest-energy. The configuration of lowest energy will be the one with the maximum number of remaining atom-atom bonds. This is an instance of the densest k-subgraph problem, with many applications beyond configurational analysis [65]. In this work, we consider only the case of removing $3$ atoms, and thus k=N3k=N-3.
@@
  • We explore the scaling behaviour of the algorithms, finding an efficient classical solution with simulated annealing that scales polynomially with system size $N$. Using simulated annealing, we solved problems with up to 388 variables. We also present some provisional evidence of quantum annealing scaling in polynomial time on the problem sizes that were feasible to run on hardware.
@@ II Related Work
More specifically, Ref. [28] demonstrates how QAA can be applied to solve a version of the graphene defect use-case formulated as a Densest k-Subgraph (DkS) problem (see Appendix A for more detail on its formulation and complexity). This problem is NP-hard for general $k$ and bipartite graphs of degree up to $3$ [67]. In [28], an $18$ atom lattice is encoded as a fully-connected QUBO and executed on quantum annealers. In our work we execute this on larger lattice sizes, using quantum annealers as well as gate-based quantum computers and classical methods. Another related piece of work [69] studies a slightly different formulation of the problem where the goal is to minimise the number of dangling bonds (i.e. connections between vacancies and carbon atoms in the lattice), again executed on a quantum annealer. The same DkS problem as in [28] has also been investigated in a biology-inspired test-case [15] on randomised bipartite graphs, where the topology of the D-Wave device is identified as the main limitation for applying it on quantum hardware.
@@ III Problem Formulation
We use the problem of finding the minimum energy of defective graphene structures as an illustrative example of our benchmarking strategy. This problem is discussed in more detail in Appendix A. We model the system as an $N$-site hexagonal structure, as shown visually in Figure 1 ($N=18$ in this example). The first graphene structure explored is a $3\times 3$ supercell, a unit cell containing two carbon atoms repeated in a $3\times 3$ arrangement, which corresponds to an $18$ variable QUBO. We also investigate other $n\times n$ supercells, which translate to QUBOs of varying size. Each (binary) QUBO variable represents a site containing either a carbon atom or a vacancy.
@@
The optimisation problem is, therefore, translated to finding the lowest energy configuration of the system, subject to the constraint of a certain number of vacancies. We work with three vacancies, the highest studied in Ref. [28], treating this number as a constant throughout our work. Holding the number of vacancies constant at three means the space of feasible solutions grows as $N\choose 3$}, where $N$ is the number of QUBO variables, which results in polynomial growth. Choosing the number of vacancies to grow as a function of $N$ would cause the feasible search space to grow at a larger rate. We experimented with this and found that it made no significant difference, for this particular QUBO instance, to the time scaling of simulated annealing. Therefore, for simplicity we kept the number of vacancies constant at three.
@@
where $\bm{x}$ is an $N$-component vector containing the binary QUBO variables that represent the different configurations, and $Q$ is the cost matrix encoding the cost function of the particular problem instance. We aim to solve the minimization problem:
@@
In order to map our problem to the above QUBO, we represent the presence of a vacancy or an atom in each site $i$ by $x_{i}$ being 0 or $1$ accordingly. The hexagonal graphene structure (with boundary conditions) can be represented by a graph with adjacency matrix $A$ and dimension $N\times N$. Since our objective is to find the maximum number of carbon-carbon bonds left after the atoms are replaced by vacancies, we want to reward (by decreasing the energy of the configuration) only the edges with a $1$ in both of their vertices. Therefore, our goal is achieved by choosing a cost matrix that is strictly upper diagonal and contains $\kappa A_{ij}$ in its non-zero elements. Parameter $\kappa$ is the bond energy, such that adding a single bond lowers the system energy by $\kappa$.
@@ IV.1 Algorithms and Parameters
The final classical method we consider is random sampling. It randomly selects a particular binary vector out of all possible bit strings $\bm{x}$ of $N$ bits (like in brute force, we do not exclude from our search space the solutions that have an incorrect number of vacancies) and evaluates its energy with respect to the cost function. This means that each binary vector has probability $1/2^{N}$ of being sampled. This is, of course, not uniform with respect to energy as multiple bitstrings can have the same energy. Random sampling provides a useful baseline for solving QUBOs with probabilistic methods, as it uses no technique that exploits the problem structure to improve the probability of finding good solutions.
@@ IV.2 Post-Selection
To perform post-selection for our use-case with a single constraint, a constant number of bitstrings must be checked, equal to the number of samples (shots), which have $N$ entries each. Therefore, the scaling of this method is linear in $N$. The time required for post-selection at the $N=18$ QUBO problem size, was found to be approximately $0.01\,\mathrm{s}$ for $10,000$ samples, which is the maximum number of samples produced as output in any of our experiments. However, for larger problems, the number of samples may need to scale with $N$ to achieve non-zero probability of the optimal solution. As $N$ increases, this probability decreases due to increased noise from using more qubits, which we have confirmed experimentally.
@@ V Results
For VQE on the state vector simulator and quantum annealing, the experiment was carried out $10$ times, the output samples were accumulated and the distributions renormalised. For VQE on the IBM QPU the experiment was carried out $5$ times due to time constraints. We report our metrics averaged across repeat experiments and their error bars as described in the previous section. Finally, as brute force is an exact method, we take it to have idealised values of our quality of solution metrics, a value of $1$ for $P_{s}$ and AR. Figure 4 shows the post-selected distributions, in which solutions that violate the constraint on the number of vacancies are excluded. We see that the distributions have more weight over a range of energies close to the minimum.
@@ V.1.1 Scaling to larger problem sizes: Quantum Annealing
We now discuss how quantum annealing runtime scales as a function of problem size. The time spent by quantum annealing on each problem instance is partly determined by the number of anneals (shots) and the annealing time. The number of anneals is chosen so that our results are statistically significant, and the annealing time is found through hyperparameter optimisation. Figure 5 displays the user runtime for simulated annealing and quantum annealing (with two different embedding techniques) as a function of the number of QUBO variables, $N$. The quantum annealing embedding technique used for the results in the previous section was minor embedding, D-Wave’s default mapping scheme, which is a heuristic method. An alternative embedding technique, ‘clique embedding’ [103], can be used for QUBO problems where the QUBO graph forms a ‘clique’, which means that it has full connectivity (100% density). Clique embedding is an algorithm that runs in polynomial time in the worst case[16, 66, 103], which has uniform chain lengths (or close to uniform) [16]. We observe near-constant scaling with clique embedding in Figure 18a. At larger problem sizes this may start to grow polynomially.
@@ Appendix A Graphene Defect Problem and its Computational Complexity
Our configurational analysis use-case is finding the energies of defective graphene structures. More specifically, given a graphene sheet which we represent by a hexagonal carbon lattice of $N$ atoms with boundary conditions, we remove some of the carbon atoms, creating vacancies on the corresponding sites of the lattice. This results in breaking of carbon-carbon bonds, or simply ‘bonds’ hereafter, between the atom we removed and its neighbours in the lattice. The energy of the resulting structure can be determined by the atoms left and their connectivity. Different approaches in determining this energy function and thus the corresponding optimisation problem have been taken in [69] and in [28].
We use the same approach as in [28]. Here the problem is the following: If we are asked to remove a specific number of atoms (i.e. create a fixed number of vacancies), the goal is to find which atoms, when removed, cause the maximum number of bonds remaining in the structure. We can model the above as a graph theoretical problem on a graph $G(V,E)$, with $|V|=N$. Each vertex $v\in V$ represents a site, which can contain an atom, in which case $v\in A$, or a vacancy, in which case $v\in\overline{A}$, where $\A\overline{A}=V\backslash$ A. The edges represent the hexagonal carbon lattice of graphene. The condition of our problem is that the size of $A$, i.e. the number of vacancies, needs to be fixed, |A|=Nk|A|=N-k. Then the problem is to determine the set $A$, i.e. the position of the vacancies, so that the bonds, which are the edges in $G$ that connect the vertices of $\overline{A}$, are maximised.
This a very common problem in complexity theory called the Densest $k$ subgraph (DkS) [65]. Notice that the configuration space, i.e. the number of possible selections of the vertices of the set $A$ of fixed size NkN-k out of $N$ sites scales with $N\choose N-k$}. For constant NkN-k (or $k$) the configuration space is polynomial on $N$ and therefore exhaustive search gives a polynomial time algorithm (note: exhaustive search for the QUBO formulation of the problem is exponential in time as the search space is $2^{N}$). For arbitrary $k$ and for general graphs the problem is known to be NP-Hard, even for graphs of degree $3$ as ours [67]. The proof is by simply observing that that the maximum clique problem can be reduced to DkS. The best known polynomial approximation algorithm has approximation ratio $N^{1/4+\epsilon}$ [26]. In the special case of bipartite graphs, such as our hexagonal carbon lattice with boundary conditions, it has been proven that the DkS problem remains hard [17]. It is also known that densest problems (more edges in the original graph $G$) are harder in general [15].
@@
A different approach in modeling the problem is taken in [69], which requires us to double the number of qubits. Here only the so-called ‘dangling bonds’, which an atom is connected to a vacancy, increase the energy of the structure and therefore its instability. Now, if we are asked to remove a specific number of atoms (a condition not enforced by the authors of [69]), the goal is to find which atoms, when removed, cause the minimum number of dangling bonds. The graph theoretical problem is therefore slightly different. Again we have a graph $G(V,E)$, with $|V|=N$ with vertices representing sites and the edges the carbon lattice, and again $A$ represents the set of vacancies. Again the problem is to find a partition of $V$ into two fixed sized sets $A$ and $\overline{A}$, but the goal now is to minimise the number of elements $e\in E$ for which $e\cap A\neq\emptyset$ and $e\cap\overline{A}\neq\emptyset$ (in other words the edges that connect one member of $A$ with one member of $\overline{A}$, are minimized). This is the well-studied minimum bisection problem (MBP), known to be NP-Hard for general graphs [87, 101]. A typical formulation of the problem is to choose $k=\lfloor N/2\rfloor$. In the special case of 33-regular graphs, such as our hexagonal carbon lattice with boundary conditions, it has been proven that the MBP problem is hard to even approximate [11]. We think that this is a very promising version of the problem to be considered in future work.
@@ D.1 RealAmplitudes vs QAOA Ansatz (Hamiltonian agnostic vs informed)
When comparing different ansatz in our hyperparameter search we found that the RealAmplitudes ansatz returned higher values of $P_{s}$ (results in Figure 8e). The simple explanation for this is that the RealAmplitudes ansatz has more parameters and is therefore more expressive allowing it to find a state with good overlap with the ground state. But what happens if we increase the depth of QAOA so that it has the same number of parameters as the RealAmplitudes ansatz? At $p=2$, for the $18$ variable problem size, QAOA has just $4$ independent parameters, but already has a depth of $58$ gates, whereas the RealAmplitudes depth is $20$ gates (these are the abstract circuits before transpilation). This is due to the fact that our problem Hamiltonian $H_{C}$, is fully connected. Consequently, the QAOA circuit has full connectivity (a fully connected entanglement arrangement), implementation of which on the IBM QPU is possible only with the introduction of SWAP gates. Running and simulating these QAOA circuits at $p=2$ can already take a long amount of time which makes increasing $p$ quite infeasible.

### v5

@@ I Introduction
To perform our cross-platform benchmarking we chose the problem of finding the lowest-energy configuration of defective graphene structures, based on the formulation found in Ref. [28]. We use the graphene defect problem solely as a representative, high density instance on which to test our benchmarking framework. Suppose we are tasked with removing a specific number of atoms from a graphene sheet of NN sites. The goal is to find which atoms, when removed, will result in a configuration with the lowest-energy. The configuration of lowest energy will be the one with the maximum number of remaining atom-atom bonds. This is an instance of the densest k-subgraph problem, with many applications beyond configurational analysis [65]. In this work, we consider only the case of removing 33 atoms, and thus k=N3k=N-3.
@@
  • We explore the scaling behaviour of the algorithms, finding an efficient classical solution with simulated annealing that scales polynomially with system size NN. Using simulated annealing, we solved problems with up to 388 variables. We also present some provisional evidence of quantum annealing scaling in polynomial time on the problem sizes that were feasible to run on hardware.
@@ II Related Work
More specifically, Ref. [28] demonstrates how QAA can be applied to solve a version of the graphene defect use-case formulated as a Densest k-Subgraph (DkS) problem (see Appendix A for more detail on its formulation and complexity). This problem is NP-hard for general $k$ and bipartite graphs of degree up to 33 [67]. In [28], an $18$ atom lattice is encoded as a fully-connected QUBO and executed on quantum annealers. In our work we execute this on larger lattice sizes, using quantum annealers as well as gate-based quantum computers and classical methods. Another related piece of work [69] studies a slightly different formulation of the problem where the goal is to minimise the number of dangling bonds (i.e. connections between vacancies and carbon atoms in the lattice), again executed on a quantum annealer. The same DkS problem as in [28] has also been investigated in a biology-inspired test-case [15] on randomised bipartite graphs, where the topology of the D-Wave device is identified as the main limitation for applying it on quantum hardware.
@@ III Problem Formulation
We use the problem of finding the minimum energy of defective graphene structures as an illustrative example of our benchmarking strategy. This problem is discussed in more detail in Appendix A. We model the system as an NN-site hexagonal structure, as shown visually in Figure 1 ($N=18$ in this example). The first graphene structure explored is a $3\times 3$ supercell, a unit cell containing two carbon atoms repeated in a $3\times 3$ arrangement, which corresponds to an $18$ variable QUBO. We also investigate other $n\times n$ supercells, which translate to QUBOs of varying size. Each (binary) QUBO variable represents a site containing either a carbon atom or a vacancy.
@@
The optimisation problem is, therefore, translated to finding the lowest energy configuration of the system, subject to the constraint of a certain number of vacancies. We work with three vacancies, the highest studied in Ref. [28], treating this number as a constant throughout our work. Holding the number of vacancies constant at three means the space of feasible solutions grows as $N\choose 3$}, where NN is the number of QUBO variables, which results in polynomial growth. Choosing the number of vacancies to grow as a function of NN would cause the feasible search space to grow at a larger rate. We experimented with this and found that it made no significant difference, for this particular QUBO instance, to the time scaling of simulated annealing. Therefore, for simplicity we kept the number of vacancies constant at three.
@@
where $\bm{x}$ is an NN-component vector containing the binary QUBO variables that represent the different configurations, and $Q$ is the cost matrix encoding the cost function of the particular problem instance. We aim to solve the minimization problem:
@@
In order to map our problem to the above QUBO, we represent the presence of a vacancy or an atom in each site $i$ by $x_{i}$ being 0 or 11 accordingly. The hexagonal graphene structure (with boundary conditions) can be represented by a graph with adjacency matrix AA and dimension $N\times N$. Since our objective is to find the maximum number of carbon-carbon bonds left after the atoms are replaced by vacancies, we want to reward (by decreasing the energy of the configuration) only the edges with a 11 in both of their vertices. Therefore, our goal is achieved by choosing a cost matrix that is strictly upper diagonal and contains $\kappa A_{ij}$ in its non-zero elements. Parameter $\kappa$ is the bond energy, such that adding a single bond lowers the system energy by $\kappa$.
@@ IV.1 Algorithms and Parameters
The final classical method we consider is random sampling. It randomly selects a particular binary vector out of all possible bit strings $\bm{x}$ of NN bits (like in brute force, we do not exclude from our search space the solutions that have an incorrect number of vacancies) and evaluates its energy with respect to the cost function. This means that each binary vector has probability $1/2^{N}$ of being sampled. This is, of course, not uniform with respect to energy as multiple bitstrings can have the same energy. Random sampling provides a useful baseline for solving QUBOs with probabilistic methods, as it uses no technique that exploits the problem structure to improve the probability of finding good solutions.
@@ IV.2 Post-Selection
To perform post-selection for our use-case with a single constraint, a constant number of bitstrings must be checked, equal to the number of samples (shots), which have NN entries each. Therefore, the scaling of this method is linear in NN. The time required for post-selection at the $N=18$ QUBO problem size, was found to be approximately $0.01\,\mathrm{s}$ for $10,000$ samples, which is the maximum number of samples produced as output in any of our experiments. However, for larger problems, the number of samples may need to scale with NN to achieve non-zero probability of the optimal solution. As NN increases, this probability decreases due to increased noise from using more qubits, which we have confirmed experimentally.
@@ V Results
For VQE on the state vector simulator and quantum annealing, the experiment was carried out $10$ times, the output samples were accumulated and the distributions renormalised. For VQE on the IBM QPU the experiment was carried out 55 times due to time constraints. We report our metrics averaged across repeat experiments and their error bars as described in the previous section. Finally, as brute force is an exact method, we take it to have idealised values of our quality of solution metrics, a value of 11 for $P_{s}$ and AR. Figure 4 shows the post-selected distributions, in which solutions that violate the constraint on the number of vacancies are excluded. We see that the distributions have more weight over a range of energies close to the minimum.
@@ V.1.1 Scaling to larger problem sizes: Quantum Annealing
We now discuss how quantum annealing runtime scales as a function of problem size. The time spent by quantum annealing on each problem instance is partly determined by the number of anneals (shots) and the annealing time. The number of anneals is chosen so that our results are statistically significant, and the annealing time is found through hyperparameter optimisation. Figure 5 displays the user runtime for simulated annealing and quantum annealing (with two different embedding techniques) as a function of the number of QUBO variables, NN. The quantum annealing embedding technique used for the results in the previous section was minor embedding, D-Wave’s default mapping scheme, which is a heuristic method. An alternative embedding technique, ‘clique embedding’ [103], can be used for QUBO problems where the QUBO graph forms a ‘clique’, which means that it has full connectivity (100% density). Clique embedding is an algorithm that runs in polynomial time in the worst case[16, 66, 103], which has uniform chain lengths (or close to uniform) [16]. We observe near-constant scaling with clique embedding in Figure 18a. At larger problem sizes this may start to grow polynomially.
@@ Appendix A Graphene Defect Problem and its Computational Complexity
Our configurational analysis use-case is finding the energies of defective graphene structures. More specifically, given a graphene sheet which we represent by a hexagonal carbon lattice of NN atoms with boundary conditions, we remove some of the carbon atoms, creating vacancies on the corresponding sites of the lattice. This results in breaking of carbon-carbon bonds, or simply ‘bonds’ hereafter, between the atom we removed and its neighbours in the lattice. The energy of the resulting structure can be determined by the atoms left and their connectivity. Different approaches in determining this energy function and thus the corresponding optimisation problem have been taken in [69] and in [28].
We use the same approach as in [28]. Here the problem is the following: If we are asked to remove a specific number of atoms (i.e. create a fixed number of vacancies), the goal is to find which atoms, when removed, cause the maximum number of bonds remaining in the structure. We can model the above as a graph theoretical problem on a graph $G(V,E)$, with $|V|=N$. Each vertex $v\in V$ represents a site, which can contain an atom, in which case $v\in A$, or a vacancy, in which case $v\in\overline{A}$, where $\A\overline{A}=V\backslash$ A. The edges represent the hexagonal carbon lattice of graphene. The condition of our problem is that the size of AA, i.e. the number of vacancies, needs to be fixed, |A|=Nk|A|=N-k. Then the problem is to determine the set AA, i.e. the position of the vacancies, so that the bonds, which are the edges in $G$ that connect the vertices of $\overline{A}$, are maximised.
This a very common problem in complexity theory called the Densest $k$ subgraph (DkS) [65]. Notice that the configuration space, i.e. the number of possible selections of the vertices of the set AA of fixed size NkN-k out of NN sites scales with $N\choose N-k$}. For constant NkN-k (or $k$) the configuration space is polynomial on NN and therefore exhaustive search gives a polynomial time algorithm (note: exhaustive search for the QUBO formulation of the problem is exponential in time as the search space is $2^{N}$). For arbitrary $k$ and for general graphs the problem is known to be NP-Hard, even for graphs of degree 33 as ours [67]. The proof is by simply observing that that the maximum clique problem can be reduced to DkS. The best known polynomial approximation algorithm has approximation ratio $N^{1/4+\epsilon}$ [26]. In the special case of bipartite graphs, such as our hexagonal carbon lattice with boundary conditions, it has been proven that the DkS problem remains hard [17]. It is also known that densest problems (more edges in the original graph $G$) are harder in general [15].
@@
A different approach in modeling the problem is taken in [69], which requires us to double the number of qubits. Here only the so-called ‘dangling bonds’, which an atom is connected to a vacancy, increase the energy of the structure and therefore its instability. Now, if we are asked to remove a specific number of atoms (a condition not enforced by the authors of [69]), the goal is to find which atoms, when removed, cause the minimum number of dangling bonds. The graph theoretical problem is therefore slightly different. Again we have a graph $G(V,E)$, with $|V|=N$ with vertices representing sites and the edges the carbon lattice, and again AA represents the set of vacancies. Again the problem is to find a partition of $V$ into two fixed sized sets AA and $\overline{A}$, but the goal now is to minimise the number of elements $e\in E$ for which $e\cap A\neq\emptyset$ and $e\cap\overline{A}\neq\emptyset$ (in other words the edges that connect one member of AA with one member of $\overline{A}$, are minimized). This is the well-studied minimum bisection problem (MBP), known to be NP-Hard for general graphs [87, 101]. A typical formulation of the problem is to choose $k=\lfloor N/2\rfloor$. In the special case of 33-regular graphs, such as our hexagonal carbon lattice with boundary conditions, it has been proven that the MBP problem is hard to even approximate [11]. We think that this is a very promising version of the problem to be considered in future work.
@@ D.1 RealAmplitudes vs QAOA Ansatz (Hamiltonian agnostic vs informed)
When comparing different ansatz in our hyperparameter search we found that the RealAmplitudes ansatz returned higher values of $P_{s}$ (results in Figure 8e). The simple explanation for this is that the RealAmplitudes ansatz has more parameters and is therefore more expressive allowing it to find a state with good overlap with the ground state. But what happens if we increase the depth of QAOA so that it has the same number of parameters as the RealAmplitudes ansatz? At $p=2$, for the $18$ variable problem size, QAOA has just 44 independent parameters, but already has a depth of $58$ gates, whereas the RealAmplitudes depth is 2020 gates (these are the abstract circuits before transpilation). This is due to the fact that our problem Hamiltonian $H_{C}$, is fully connected. Consequently, the QAOA circuit has full connectivity (a fully connected entanglement arrangement), implementation of which on the IBM QPU is possible only with the introduction of SWAP gates. Running and simulating these QAOA circuits at $p=2$ can already take a long amount of time which makes increasing $p$ quite infeasible.